# HCF: Hierarchical Cascade Framework for Distributed Multi-Stage Image Compression

### v2

@@ 1 Introduction
The rapid growth of digital media consumption has significantly increased demands on visual information transmission systems (GHOSH2025108011; 6G_Visual_Data), positioning image compression as an essential technology to address bandwidth and storage constraints while maintaining acceptable quality levels (ibraheem2024comprehensive). Modern transmission scenarios involve multi-stage processing where visual content traverses multiple processing nodes during transmission (10.1145/3742472). As deployment scenarios become increasingly complex, it is imperative for compression systems to provide multiple quality levels, enabling quality scaling during transmission (kong2024progressive). This creates scenarios where compression operations are distributed across multiple processing stages with varying quality levels—a paradigm we refer to as distributed multi-stage image compression.
Within the distributed multi-stage compression framework, existing approaches have emerged but face fundamental limitations. Progressive compression methods (Lee_2022_CVPR; 2023_CVPR_jeon; lee2025deephq; Presta_2025_WACV) enable quality adaptation through bitstream truncation but constrain intermediate processing to passive operations with suboptimal rate-distortion trade-offs. Successive compression approaches (SIC) achieve adaptation through repeated operations but suffer from cumulative quality degradation and computational inefficiency due to redundant pixel-domain conversions. Meanwhile, fixed-parameter compression models (minnenbt18; cheng2020image; jiang2023mlicpp), while achieving superior rate-distortion performance in centralized settings, provide single operating points without post-encoding flexibility required for distributed multi-stage scenarios.
The emergence of distributed computational resources (MobileEdgeComputing2025; 6g_edge; TangComputingNetworkConvergence; INCML; INC2023) in modern networks creates new opportunities for intelligent quality transformation within distributed multi-stage compression, necessitating methods that can efficiently leverage these capabilities while avoiding the performance penalties of existing approaches. To address this need, we propose the Hierarchical Cascade Framework (HCF), which harnesses distributed computational resources for efficient multi-stage image compression. To the best of our knowledge, this is the first framework to effectively leverage distributed computational resources for multi-stage compression by enabling direct latent-space transformations across network nodes and implementing policy-driven quantization control. This approach simultaneously improves compression performance and reduces computational overhead while preventing cumulative degradation through optimal placement strategies.
@@ 2 Related Work
Image compression has evolved from traditional standards (jpeg; jpeg2000; hevc; webp; bpg) to deep learning approaches that achieve superior rate-distortion performance. This paradigm shift was marked by key milestones: autoencoder architectures with neural transforms (balle2016end; Lossy_2017), entropy modeling through hyperprior networks (ballemshj18), scale-mean Gaussians (minnenbt18), and mixture models (cheng2020image; zhu2022unified; fu2023), alongside context-aware processing that exploited local spatial dependencies (minnenbt18; he2022; liu2023learned; liu2024; jiang2023mlicpp), global correlations (guo2021causal; khoshkhahtinat2023multi), and channel-wise relationships (feng2025linear; zeng2025mambaic), as well as dictionary-based modeling (wu2025conditional; lu2025learned) that leverages external priors. However, these learned methods remain designed as single-stage, centralized systems that generate static bitstreams, limiting their optimization potential in scenarios requiring post-encoding quality adaptation.
@@
To address quality adaptation requirements, progressive compression approaches have evolved from early foundations with recurrent neural networks for variable compression rates (toderici2015variable; Toderici) and spatial adaptive bit allocation and diffusion mechanisms (Johnston). These advances enabled scalable coding architectures with layer-wise bit allocation (su), and fine granular scalability through nested quantization (lu2021progressive), trit-plane encoding (Lee_2022_CVPR; 2023_CVPR_jeon), spatial autoregressive modeling with codeword alignment (Tian_VCIP), multirate progressive entropy modeling that unifies split-coded-then-merge models (Li_tcsvt), progressive deep coding with layer-specific parameter learning (lee2025deephq) and element-wise progressive transmission via variance-aware masking (Presta_2025_WACV).
@@
Meanwhile, successive image compression pipelines, where the same codec is repeatedly applied across stages via encoding–decoding cycles, can in principle produce multiple quality levels, but have been shown to suffer from cumulative quality degradation and instability (SIC). In parallel, the emergence of distributed computational resources in modern networks—including 6G Edge Networks (6g_edge), Mobile Edge Computing (choi2025cooperative; MobileEdgeComputing2025), and In-Network Machine Learning (INCML)—highlights a broader shift towards exploiting intermediate nodes as active compute and decision points. These capabilities provide new opportunities for compression optimization through intermediate computational processing (10.1145/3742472). However, existing compression methods cannot effectively leverage such resources: progressive methods are constrained to suboptimal truncation operations with passive intermediate nodes, whereas successive approaches incur computational redundancy through repeated pixel-domain operations.
@@ 3.1 Existing Frameworks and Limitations
Distributed Recompression Framework (DRF). DRF extends Successive Image Compression (SIC) (SIC) to distributed multi-stage scenarios. Let $s$ denote the source quality level, $d$ the target quality level, with $s\geq d$, and let $k\in\{s,s-1,\ldots,d\}$ denote the stage index; a lower index implies a more compressed state. Under DRF, as shown in Figure 1, each intermediate node performs quality adaptation through pixel-level recompression: receiving compressed data at quality level sk+1s-k+1, applying decompression operations $D^{s-k+1},g_{s}^{s-k+1}$, then compression operations $g_{a}^{s-k},Q^{s-k},E^{s-k}$ to achieve quality level sks-k. Although it actively utilizes intermediate computational resources, it suffers from redundant pixel-domain conversions and computational inefficiency.
@@ 3.2 Hierarchical Cascade Framework
Note that both the inter- and intra-node processes end with a transform module that shares an identical architecture (see Figure 2(b)) but requires separate training for different input distributions at each level. The novel modules contain residual blocks, attention mechanisms, convolution operations with GDN (GDN) and LeakyReLU activations, and channel dimension adjustments to handle both quantized and unquantized latent representations. Our primary contribution lies in the development of a systematic framework enabling policy-driven quantization control across multiple compression stages, with the unified architecture facilitating seamless integration across the compression pipeline. Ablation studies (Section 4.5) demonstrate the effectiveness of this design and confirm superior performance in distributed multi-stage scenarios.
@@ 3.3 Training Strategy
where $\tilde{y}^{k}=(\mathcal{F}_{s\rightarrow k}^{\boldsymbol{\pi}}\circ g_{a}^{s})(x)$ and $\hat{y}^{k}=D^{k}\!\big(E^{k}\!\big(Q^{k}(\tilde{y}^{k})\big)\big)$. During training, all previously optimized networks are frozen, and only the current transform module is updated. Our unified architecture handles both quantized and unquantized inputs via weight switching rather than separate models, minimizing deployment overhead. Inspired by knowledge distillation (hinton2015distilling), our modules enable adaptive compression across distributed quality levels in multi-stage compression systems.
@@ 4.1 Experimental Setup
We validated HCF across five diverse compression architectures—MLIC++ (jiang2023mlicpp), cheng2020_attn and cheng2020_anchor (cheng2020image), and mbt2018 and mbt2018_mean (minnenbt18)—spanning from hyperprior models to context-attention mechanisms, ensuring generalization across architectural paradigms. We train the models following our two-phase strategy (Section 3.3) using DUTS dataset (dataset) with standard augmentation. The models are initialized from pre-trained CompressAI (compressai) weights and official implementations with quality levels $s=6$ (MLIC++, cheng2020_anchor, cheng2020_attn) and $s=8$ (mbt2018, mbt2018_mean). We use Kodak (kodak_dataset), CLIC2020-mobile (clic2020), and CLIC (clic2020) datasets for the evaluation.
We systematically compared HCF against four baseline categories: (1) PCF including latest SOTA approaches (Presta (Presta_2025_WACV), DeepHQ (lee2025deephq), Li (Li_tcsvt), Jeon (2023_CVPR_jeon), Tian (Tian_VCIP), Lee (Lee_2022_CVPR)) and foundational works (Toderici; Johnston; Diao); (2) DRF extending successive compression (SIC); (3) SSF as performance bounds; (4) Traditional standards (JPEG (jpeg), JPEG2000 (jpeg2000), BPG (bpg), HEVC (hevc), WebP (webp)).
For a fair comparison, DRF and SSF use identical architectures to HCF. We report rate–distortion curves and Bjøntegaard Delta metrics (BJONTEGARD): $_{\text{P}}/BD-Rate$$_{\text{M}}$ (bitrate reduction for PSNR/MS-SSIM (Wang_msssim), lower is better) and BD-PSNR/BD-MS-SSIM (quality improvement, higher is better), along with computational efficiency metrics including FLOPs, parameters, GPU memory, and execution time reduction relative to the baselines. All experiments used NVIDIA RTX A6000 GPUs. The experimental details and hyperparameters are described in Supplementary Material.
@@ 4.2 Policy Analysis
To understand the underlying mechanism driving this performance difference, we analyzed differential entropy evolution across compression levels using the Kozachenko-Leonenko estimator (KozLeo87). Figure 3 shows differential entropy evolution across four compression levels for three quantization policies. While all policies converge to similar entropy values before the final quantization operation, they exhibit markedly different entropy increases after quantization: 94.1, 96.1, and 99.3 bits for policies [1,0,0,1], [0,1,0,1], and [0,0,1,1], respectively. Policy $\boldsymbol{\pi}^{\text{edge}}=[1,0,0,1]$ achieved the smallest increment, demonstrating that early quantization injection allows subsequent transform modules to more effectively decorrelate and suppress redundant information, thereby minimizing the irreducible uncertainty that must be encoded. This theoretical advantage translated into visually perceptible improvements, as demonstrated in Figure 4, where $\boldsymbol{\pi}^{\text{edge}}$ preserves significantly more structural and textural details compared to alternative policies at equivalent bitrates. Comprehensive entropy analysis across different quality configurations (detailed in Supplementary Material) confirms that this pattern consistently holds across diverse scenarios.
@@ 4.3 Performance Comparison
Figure 5: Rate-distortion comparison on Kodak (PSNR) and CLIC (MS-SSIM) datasets. HCF with optimal $\boldsymbol{\pi}^{\text{edge}}$ policy consistently outperforms distributed baselines while approaching centralized SSF performance. Shaded areas indicate HCF’s superior performance over DRF.
@@
Table 3: Computational efficiency comparison: HCF vs. DRF across five multi-stage compression nodes. Values show HCF’s percentage reduction in FLOPs/Parameters/GPU Memory/Execution Time relative to DRF. $\uparrow$ higher is better.
@@ 4.5 Ablation Studies
To validate our design hypothesis that both transform modules are essential and complementary, we compared three HCF configurations: complete HCF ($\phi^{\text{inter}}+\phi^{\text{intra}}$), $\phi^{\text{inter}}$-only, and $\phi^{\text{intra}}$-only. All use $\boldsymbol{\pi}^{\text{edge}}$ policy. Figure 6 shows that removing either module leads to performance degradation: $\phi^{\text{inter}}$-only suffered at lower bitrates due to inadequate information preservation, while $\phi^{\text{intra}}$-only fails at higher bitrates due to poor quantization artifact handling. These results confirmed that both modules address distinct challenges and their combination is essential for optimal performance, with additional ablation studies in Supplementary.
Figure 6: Ablation study comparing complete HCF with single-module variants on CLIC2020-mobile dataset. Complete HCF consistently outperforms both alternatives.

### v3

@@ 1 Introduction
The rapid growth of digital media consumption has significantly increased demands on visual information transmission systems (Ghosh and Singhal 2025; Cai et al. 2024), positioning image compression as an essential technology to address bandwidth and storage constraints while maintaining acceptable quality levels (Ibraheem et al. 2024). Modern transmission scenarios involve multi-stage processing where visual content traverses multiple processing nodes during transmission (Peroni and Gorinsky 2025). As deployment scenarios become increasingly complex, it is imperative for compression systems to provide multiple quality levels, enabling quality scaling during transmission (Kong and Sun 2024). This creates scenarios where compression operations are distributed across multiple processing stages with varying quality levels—a paradigm we refer to as distributed multi-stage image compression.
Within the distributed multi-stage compression framework, existing approaches have emerged but face fundamental limitations. Progressive compression methods (Lee et al. 2022; Jeon et al. 2023; Lee et al. 2025; Presta et al. 2025) enable quality adaptation through bitstream truncation but constrain intermediate processing to passive operations with suboptimal rate-distortion trade-offs. Successive compression approaches (Kim et al. 2022) achieve adaptation through repeated operations but suffer from cumulative quality degradation and computational inefficiency due to redundant pixel-domain conversions. Meanwhile, fixed-parameter compression models (Minnen et al. 2018; Cheng et al. 2020; Jiang et al. 2025), while achieving superior rate-distortion performance in centralized settings, provide single operating points without post-encoding flexibility required for distributed multi-stage scenarios.
The emergence of distributed computational resources (Yang et al. 2025; Lin et al. 2024; Tang et al. 2024; Zheng et al. 2024; Kianpisheh and Taleb 2023) in modern networks creates new opportunities for intelligent quality transformation within distributed multi-stage compression, necessitating methods that can efficiently leverage these capabilities while avoiding the performance penalties of existing approaches. To address this need, we propose the Hierarchical Cascade Framework (HCF), which harnesses distributed computational resources for efficient multi-stage image compression. To the best of our knowledge, this is the first framework to effectively leverage distributed computational resources for multi-stage compression by enabling direct latent-space transformations across network nodes and implementing policy-driven quantization control. This approach simultaneously improves compression performance and reduces computational overhead while preventing cumulative degradation through optimal placement strategies.
@@ 2 Related Work
Image compression has evolved from traditional standards (Wallace 1992; Skodras et al. 2001; Sullivan et al. 2012; Lian and Shilei 2012; Bellard 2014) to deep learning approaches that achieve superior rate-distortion performance. This paradigm shift was marked by key milestones: autoencoder architectures with neural transforms (Ballé et al. 2017; Theis et al. 2017), entropy modeling through hyperprior networks (Ballé et al. 2018), scale-mean Gaussians (Minnen et al. 2018), and mixture models (Cheng et al. 2020; Zhu et al. 2022; Fu et al. 2023), alongside context-aware processing that exploited local spatial dependencies (Minnen et al. 2018; He et al. 2022; Liu et al. 2023, 2024; Jiang et al. 2025), global correlations (Guo et al. 2021; Khoshkhahtinat et al. 2023), and channel-wise relationships (Feng et al. 2025; Zeng et al. 2025), as well as dictionary-based modeling (Wu et al. 2025; Lu et al. 2025) that leverages external priors. However, these learned methods remain designed as single-stage, centralized systems that generate static bitstreams, limiting their optimization potential in scenarios requiring post-encoding quality adaptation.
@@
To address quality adaptation requirements, progressive compression approaches have evolved from early foundations with recurrent neural networks for variable compression rates (Toderici et al. 2016, 2017) and spatial adaptive bit allocation and diffusion mechanisms (Johnston et al. 2018). These advances enabled scalable coding architectures with layer-wise bit allocation (Su et al. 2020), and fine granular scalability through nested quantization (Lu et al. 2021), trit-plane encoding (Lee et al. 2022; Jeon et al. 2023), spatial autoregressive modeling with codeword alignment (Tian et al. 2023), multirate progressive entropy modeling that unifies split-coded-then-merge models (Li et al. 2024), progressive deep coding with layer-specific parameter learning (Lee et al. 2025) and element-wise progressive transmission via variance-aware masking (Presta et al. 2025).
@@
Meanwhile, successive image compression pipelines, where the same codec is repeatedly applied across stages via encoding–decoding cycles, can in principle produce multiple quality levels, but have been shown to suffer from cumulative quality degradation and instability (Kim et al. 2022). In parallel, the emergence of distributed computational resources in modern networks—including 6G Edge Networks (Lin et al. 2024), Mobile Edge Computing (Lee et al. 2024; Yang et al. 2025), and In-Network Machine Learning (Zheng et al. 2024)—highlights a broader shift towards exploiting intermediate nodes as active compute and decision points. These capabilities provide new opportunities for compression optimization through intermediate computational processing (Peroni and Gorinsky 2025). However, existing compression methods cannot effectively leverage such resources: progressive methods are constrained to suboptimal truncation operations with passive intermediate nodes, whereas successive approaches incur computational redundancy through repeated pixel-domain operations.
@@ 3.1 Existing Frameworks and Limitations
Distributed Recompression Framework (DRF). DRF extends Successive Image Compression (SIC) (Kim et al. 2022) to distributed multi-stage scenarios. Let $s$ denote the source quality level, $d$ the target quality level, with $s\geq d$, and let $k\in\{s,s-1,\ldots,d\}$ denote the stage index; a lower index implies a more compressed state. Under DRF, as shown in Figure 1, each intermediate node performs quality adaptation through pixel-level recompression: receiving compressed data at quality level sk+1s-k+1, applying decompression operations $D^{s-k+1},g_{s}^{s-k+1}$, then compression operations $g_{a}^{s-k},Q^{s-k},E^{s-k}$ to achieve quality level sks-k. Although it actively utilizes intermediate computational resources, it suffers from redundant pixel-domain conversions and computational inefficiency.
@@ 3.2 Hierarchical Cascade Framework
Note that both the inter- and intra-node processes end with a transform module that shares an identical architecture (see Figure 2(b)) but requires separate training for different input distributions at each level. The novel modules contain residual blocks, attention mechanisms, convolution operations with GDN (Ballé et al. 2016) and LeakyReLU activations, and channel dimension adjustments to handle both quantized and unquantized latent representations. Our primary contribution lies in the development of a systematic framework enabling policy-driven quantization control across multiple compression stages, with the unified architecture facilitating seamless integration across the compression pipeline. Ablation studies (Section 4.5) demonstrate the effectiveness of this design and confirm superior performance in distributed multi-stage scenarios.
@@ 3.3 Training Strategy
where $\tilde{y}^{k}=(\mathcal{F}_{s\rightarrow k}^{\boldsymbol{\pi}}\circ g_{a}^{s})(x)$ and $\hat{y}^{k}=D^{k}\!\big(E^{k}\!\big(Q^{k}(\tilde{y}^{k})\big)\big)$. During training, all previously optimized networks are frozen, and only the current transform module is updated. Our unified architecture handles both quantized and unquantized inputs via weight switching rather than separate models, minimizing deployment overhead. Inspired by knowledge distillation (Hinton et al. 2015), our modules enable adaptive compression across distributed quality levels in multi-stage compression systems.
@@ 4.1 Experimental Setup
We validated HCF across five diverse compression architectures—MLIC++ (Jiang et al. 2025), cheng2020_attn and cheng2020_anchor (Cheng et al. 2020), and mbt2018 and mbt2018_mean (Minnen et al. 2018)—spanning from hyperprior models to context-attention mechanisms, ensuring generalization across architectural paradigms. We train the models following our two-phase strategy (Section 3.3) using DUTS dataset (Wang et al. 2017) with standard augmentation. The models are initialized from pre-trained CompressAI (Bégaint et al. 2020) weights and official implementations with quality levels $s=6$ (MLIC++, cheng2020_anchor, cheng2020_attn) and $s=8$ (mbt2018, mbt2018_mean). We use Kodak (Kodak 1993), CLIC2020-mobile (Toderici et al. 2020), and CLIC (Toderici et al. 2020) datasets for the evaluation.
We systematically compared HCF against four baseline categories: (1) PCF including latest SOTA approaches (Presta (Presta et al. 2025), DeepHQ (Lee et al. 2025), Li (Li et al. 2024), Jeon (Jeon et al. 2023), Tian (Tian et al. 2023), Lee (Lee et al. 2022)) and foundational works (Toderici et al. 2017; Johnston et al. 2018; Diao et al. 2020); (2) DRF extending successive compression (Kim et al. 2022); (3) SSF as performance bounds; (4) Traditional standards (JPEG (Wallace 1992), JPEG2000 (Skodras et al. 2001), BPG (Bellard 2014), HEVC (Sullivan et al. 2012), WebP (Lian and Shilei 2012)).
For a fair comparison, DRF and SSF use identical architectures to HCF. We report rate–distortion curves and Bjøntegaard Delta metrics (Bjøntegaard 2001): $_{\text{P}}/BD-Rate$$_{\text{M}}$ (bitrate reduction for PSNR/MS-SSIM (Wang et al. 2003), lower is better) and BD-PSNR/BD-MS-SSIM (quality improvement, higher is better), along with computational efficiency metrics including FLOPs, parameters, GPU memory, and execution time reduction relative to the baselines. All experiments used NVIDIA RTX A6000 GPUs. The experimental details and hyperparameters are described in Supplementary Material.
@@ 4.2 Policy Analysis
To understand the underlying mechanism driving this performance difference, we analyzed differential entropy evolution across compression levels using the Kozachenko-Leonenko estimator (Kozachenko and Leonenko 1987). Figure 3 shows differential entropy evolution across four compression levels for three quantization policies. While all policies converge to similar entropy values before the final quantization operation, they exhibit markedly different entropy increases after quantization: 94.1, 96.1, and 99.3 bits for policies [1,0,0,1], [0,1,0,1], and [0,0,1,1], respectively. Policy $\boldsymbol{\pi}^{\text{edge}}=[1,0,0,1]$ achieved the smallest increment, demonstrating that early quantization injection allows subsequent transform modules to more effectively decorrelate and suppress redundant information, thereby minimizing the irreducible uncertainty that must be encoded. This theoretical advantage translated into visually perceptible improvements, as demonstrated in Figure 4, where $\boldsymbol{\pi}^{\text{edge}}$ preserves significantly more structural and textural details compared to alternative policies at equivalent bitrates. Comprehensive entropy analysis across different quality configurations (detailed in Supplementary Material) confirms that this pattern consistently holds across diverse scenarios.
@@ 4.3 Performance Comparison
Figure 5: Rate-distortion comparison using PSNR on Kodak dataset and MS-SSIM on CLIC dataset. HCF with $\boldsymbol{\pi}^{\text{edge}}$ policy outperforms PCF and DRF baselines while approaching SSF performance. Shaded areas indicate HCF’s performance advantage over DRF. Upper-left is better.
@@
Table 3: Computational efficiency comparison: HCF vs. DRF across five processing nodes. Values show HCF’s percentage reduction in FLOPs/Parameters/GPU Memory/Execution Time relative to DRF for transform operations. $\uparrow$ higher is better.
@@ 4.5 Ablation Studies
To validate our design hypothesis that both transform modules are essential and complementary, we compared three HCF configurations: complete HCF ($\phi^{\text{inter}}+\phi^{\text{intra}}$), $\phi^{\text{inter}}$-only, and $\phi^{\text{intra}}$-only. All use $\boldsymbol{\pi}^{\text{edge}}$ policy. Figure 6 shows that removing either module leads to performance degradation: $\phi^{\text{inter}}$-only suffered at lower bitrates due to inadequate information preservation, while $\phi^{\text{intra}}$-only fails at higher bitrates due to poor quantization artifact handling. These results confirmed that both modules address distinct challenges and their combination is essential for superior performance, with additional ablation studies in Supplementary.
Figure 6: Ablation study comparing complete HCF ($\phi^{\text{inter}}+\phi^{\text{intra}}$) with $\phi^{\text{inter}}$-only and $\phi^{\text{intra}}$-only variants on CLIC2020-mobile dataset under $\boldsymbol{\pi}^{\text{edge}}$ policy. Both modules are essential for superior performance. Upper-left is better.